Please help, unsure how to start this.
Answer #1

using conservation of energy

total KE at infinity = max potential energy at approach

2* (0.5* m v^2) = k q^2 / r

1.67*10^-27* (1.5*10^3)^2 = 9*10^9* (1.6*10^-19)^2 / r

r = 6.13*10^-8 m

========

Comment in case any doubt, will reply for sure.. Goodluck